# Machine learning of the Ising model on a spherical Fibonacci lattice

### v2

@@ III.3 Determining $T_{c}$ via the method of machine learning
The second approach employs machine learning, which has proven to be a powerful tool in processing big data [37, 38, 26, 39, 40, 41, 42, 43]. Given that the Fibonacci lattice is not homogeneous and does not resemble a typical image-like structure, traditional CNNs can not be applied to this model [44, 45]. Instead, we utilize GCNs, which are capable of capturing spatial structural features by leveraging the connectivity relationships between nodes [46]. This makes GCNs particularly well-suited for processing topologically structured data like the Fibonacci lattice. Details of GCN implementation can be found in Appendix C.
Figure 8: Classification confidences, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature for the 2D Ising model on a 20$\times$20 square lattice (Top panel) and 40$\times$40 square lattice (Bottom panel).
@@
To determine the critical temperature $T_{c}$ using GCNs, we adopt a supervised learning approach inspired by prior studies [33], but tailored to our system. Unlike traditional methods that rely on training samples labeled by a classifier with prior knowledge of $T_{c}$, we take advantage of the clear separation of phases at extreme temperatures to construct our training set. Specifically, Monte Carlo (MC) spin configurations are generated at very low temperatures (e.g., $T/J=0.01$, well below the expected $T_{c}$) and very high temperatures (e.g., $T/J=8.0$, well above the expected $T_{c}$), where the system is unambiguously in the ordered and disordered phases, respectively. These configurations are labeled as “ordered” ($p_{\text{o}}=1$, $p_{\text{d}}=0$) and “disordered” ($p_{\text{o}}=0$, $p_{\text{d}}=1$) based on their temperature regimes, eliminating the need for an exact $T_{c}$ value a priori. Here $p_{\text{o}}$ and $p_{\text{d}}$ represent the classification confidences for the ordered and disordered phases, respectively. The trained GCN then interpolates across the temperature range $[0.01,8.0]J$, predicting classification confidences $p_{\text{o}}$ and $p_{\text{d}}$ for each temperature. The intersection point where $p_{\text{o}}=p_{\text{d}}$ is taken as the estimated $T_{c}$.
@@
To verify the accuracy of the GCN simulations, we first investigate two-dimensional square lattices with sizes of $20\times 20$ and $40\times 40$ as a preliminary test. As illustrated in Fig.8 (with the black solid line representing $p_{\text{o}}$ and the red dashed line representing $p_{\text{d}}$), we observe that as the temperature increases, the black solid line gradually declines from 1, indicating a decreasing similarity between the simulation results and the ordered phase. In contrast, the red dashed line steadily increases from 0, reflecting an increasing similarity between the simulation results and the disordered phase. Through the simulation and analyses of different lattice scales, it is found that as the number of lattice sites increases, the phase transition temperature gradually approaches the theoretical value of $T_{c}/J=2.269$. Specifically, the values obtained are $T_{c}/J=2.51$ for the $20\times 20$ lattice and $T_{c}/J=2.36$ for the $40\times 40$ lattice. The latter demonstrates strong concordance with the predicted critical temperature. This indicates that increasing the lattice size can enhance the accuracy of phase transition temperature estimations. Moreover, the GCN effectively captures the phase transition behavior and critical properties.
@@
After validating the accuracy of GCN, we subsequently apply it to determine the phase transition temperature of the spherical Ising model. In numerical calculations, we select 800 temperature nodes within the range of $[0.01,8.0]J$, with an interval of $0.01J$, and perform 100 simulations for each temperature node. The results, presented in the top panel of Fig.9, indicate that the critical temperature is determined to be $T_{c}/J=2.35$, where $p_{\text{o}}=p_{\text{d}}$. This value is fairly close to those obtained from the analysis of the specific heat $C_{V}$, but it is comparatively higher than that of the square lattice, $T^{\Box}_{c}=2.269J$. Due to the inhomogeneity of the Fibonacci lattice, a complete analytical analysis is not feasible. We believe a possible reason for this is the presence of sites with 3 or 5 neighbors, which affects the spin dynamics within the system. At low temperatures, a site with a greater number of neighbors is less likely to experience a spin flip, as confirmed in Fig.4. However, as the temperature increases beyond $T\gtrsim 2.34J$, the situation changes. Fig.4 shows that at $T=2.34J$, the spin-flip probabilities follow the order $P_{4}>P_{5}>P_{3}$ at $T=2.34J$, indicating that spins on sites with 3 or 5 neighbors are more resistant to flipping. This, to some extent, enhances the “stability” of the ordered phase, effectively elevating the phase transition temperature.
@@
To analyze the uncertainty of $T_{c}$, we note that it is determined by the intersection of the two curves $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$, both of which exhibit good linearity near $T_{c}$. Therefore, the uncertainty can be estimated as follows. We select an interval $[T_{1},T_{2}]$ around $T_{c}$, where $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$ can be approximated as linear functions with slopes $k_{\text{o}}$ and $k_{\text{d}}$, respectively. Each point on these two curves has its own uncertainty, obtained by averaging over 10 stable spin configurations. Using error propagation, we can determine the uncertainties of $k_{\text{o}}$ and $k_{\text{d}}$, denoted as $\sigma_{\text{o}}$ and $\sigma_{\text{d}}$, respectively. Finally, the uncertainty of $T_{c}$ is given by
@@
In the bottom panel of Fig.9, we present our numerical results, yielding $T_{c}=2.338\pm 0.0784J$. Note the value $2.338J$ is obtained from the intersection of the two linear fits, which differs slightly from $T_{c}=2.35$, determined by the condition $p_{\text{o}}=p_{\text{d}}$.
III.4 Effects of $r_{c}$ and $N$
$r_{c}/R$ $N_{3}$ ($w_{3}$) $N_{4}$ ($w_{4}$) $N_{5}$ ($w_{5}$) $N_{6}$ ($w_{6}$) $N_{7}$ ($w_{7}$) $N_{8}$ ($w_{8}$)
1.298 76 (7.6%) 850 (85%) 74 (7.4%) - - -
1.500 - 26 (2.6%) 240 (24%) 734 (73.4%) - -
1.700 - - 8 (0.8%) 722 (72.2%) 224 (22.4%) 46 (4.6%)
@@ III.4 Effects of $r_{c}$ and $N$
Now, we investigate the effects of other parameters, such as $r_{c}$ and $N$, on $T_{c}$. First, we fix the total number of sites at $N=1000$ and vary the interaction length $r_{c}$. As $r_{c}$ increases, each spin interacts with a greater number of spins, effectively increasing the number of neighbors per site and altering the lattice structure. In this scenario, as the temperature rises, the energy cost for a spin flip becomes larger due to the higher average number of neighbors per site, leading to an effective increase in $T_{c}$. We present our numerical results in Fig.10. The top panel shows that $T_{c}=2.35J$ for $r_{c}=0.1298R$, $T_{c}=3.52J$ for $r_{c}=0.1500R$, and $T_{c}=3.95J$ for $r_{c}=0.1700R$, which confirms our former reasoning. Moreover, the error made by GCN when performing phase classification can serve as an indicator for the critical behavior of a dynamical system. Consequently, the temperature at which the classification error is maximized can itself be used as a phase transition criterion [47]. To quantify this, we perform 10 independent simulations at each temperature point and compute the variance of the classification confidence. As shown in the middle panel, the maximum classification errors occur at $T=2.34J$ for $r_{c}=0.1298R$, $T_{c}=3.515J$ for $r_{c}=0.1500R$, and $T_{c}=3.961J$ for $r_{c}=0.1700R$, which agree well with the phase transition temperature obtained from the intersection of the two phases ($p_{\text{o}}$, $p_{\text{d}}$) in classification. This further supports the reliability of the maximum confusion criterion. The bottom panel lists the fraction of sites with different numbers of neighbors for various $r_{c}$ values. For example, when $r_{c}=0.1500R$, $73.4\%$ of the sites have six neighbours, indicating that most of the lattice resembles a triangular structure. Interestingly, the estimated $T_{c}=3.52J$ is relatively close to the exact value $T_{c}=4/\text{ln}3\approx 3.64J$ of the planar triangular lattice as reported in Ref.[48]. Since this model also includes some four-neighbor and five-neighbor sites, its spin flip energy is lower than that of the six-neighbor sites, which results in a relatively lower $T_{c}$. For $r_{c}=0.1700R$, although a similar fraction of sites have six neighbors ($w_{6}=72.2\%$), the number of sites with more neighbors increases significantly ($w_{7}=22.4\%$ and $w_{8}=4.6\%$). This enhanced ”neighbor interaction” effectively strengthens spin correlations, leading to a higher phase transition temperature of $T_{c}=3.92J$.
$N$ $r_{c}$ $N_{3}$ ($w_{3}$) $N_{4}$ ($w_{4}$) $N_{5}$ ($w_{5}$)
500 1.8130 68 (11.6%) 418 (83.6%) 14(2.8%)
1000 1.2980 76 (7.6%) 850 (85%) 74 (7.4%)
3000 0.7470 186 (6.2%) 2740 (91.3%) 74 (2.5%)
5000 0.5818 112 (2.2%) 4740 (94.8%) 148 (3%)
@@
Figure 11: (Top panle) Classification confidences, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature for $(N,r_{c}/R)=(1000,0.1298)$, (3000,0.747) and (5000,0.5818). (Bottom panel) Corresponding fractions of sites with different numbers of neighbors.
@@
Next, we vary the number of lattice sites and adjust $r_{c}$ accordingly to ensure the majority of sites have four neighbors. However, due to the absence of translational and rotational symmetry in the spherical Fibonacci lattice, this adjustment does not necessarily guarantee that lattices with different site numbers share similar geometric structures. This contrasts with the 2D square lattice, where structural consistency is maintained. In addition to the previously discussed case of $N=1000$, $r_{c}/R=0.1298$, we also select $N=500$, $r_{c}/R=1.8130$; $N=3000$, $r_{c}/R=0.7470$ and $N=5000$, $r_{c}/R=0.5818$, ensuring that the fractions of sites with four neighbors in the latter three cases are 83.6%, 91.3% and 94.8%, respectively. Interestingly, this suggests that the “irregular area” gradually diminishes as the effective curvature of the spherical lattice decreases. The numerical results are visualized in the top panel of Fig.11. Similarly, we determine the phase transition temperatures using two different methods mentioned before. First, by identifying the intersection of the classification confidences, we obtain $T_{c}=2.31J$, $2.35J$, $2.34J$ and $2.33J$ for $N=500$, $1000$, $3000$ and $5000$, respectively. Second, using linear fitting near $T_{c}$ (see the bottom panel of Fig.9), the corresponding phase transition temperatures are giving $T_{c}=2.330\pm 0.1130J$, $2.338\pm 0.0784J$, $2.336\pm 0.0425J$ and $2.347\pm 0.0479J$ for the corresponding system sizes. Surprisingly, at least within the margin of error, these results do not reflect the impact of finite system size on the phase transition temperature. We hypothesize that the following reasons may contribute to this observation: 1. Finite-size effects may not be prominent enough. Although the number of sites increases tenfold from $500$ to $5000$ for the same spherical surface, the interaction range $r_{c}/R$ only decreases from $1.813$ to $0.5818$, a reduction by a factor of $3.12$. More pronounced finite-size effects may require considering systems with significantly larger numbers of sites, which would demand greater computational resources. 2. The folding of the lattice onto the sphere preserves uniformity, maintaining a similarity between local and global behaviors $N\to\infty$. This inherent uniformity may suppress noticeable finite-size effects. All these findings suggest that even for a simple system like the Ising model, the spherical Fibonacci lattice exhibits properties distinct from those of planar lattices, similar to the observations in Ref.[27].

### v3

@@ III.3 Determining $T_{c}$ via the method of GCNs
The second approach employs machine learning, which has proven to be a powerful tool in processing big data Wang (2016); Carleo and Troyer (2017); Carrasquilla and Melko (2017); Ch’ng et al. (2017); Deng et al. (2017); Wetzel (2017a); Zhang et al. (2018); Venderley et al. (2018). Given that the Fibonacci lattice is not homogeneous and does not resemble a typical image-like structure, traditional CNNs can not be applied to this model Beach et al. (2018); Zhang et al. (2019). Instead, we utilize GCNs, which are capable of capturing spatial structural features by leveraging the connectivity relationships between nodes Zhou et al. (2020). This makes GCNs particularly well-suited for processing topologically structured data like the Fibonacci lattice. The GCNs does not require explicit geometric coordinates of the spherical Fibonacci lattice (e.g., site latitude and longitude) and can capture local spin patterns through the adjacency matrix based on a nearest-neighbor cutoff radius $r_{s}$. Because strong correlations between spins imprint phase transition information onto local patterns, the GCNs can infer global phases from local structures. This behavior was also observed in our previous study of the spherical XY model Song et al. (2022). Details of GCNs implementation can be found in Appendix B.
Figure 7: Classification confidences $p_{\text{o}}$ and $p_{\text{d}}$ as a function of temperature for the 2D Ising model on a $L\times L$ square lattice with periodic boundary conditions. The pink, red, black, blue, and green lines correspond to lattice sizes $L=20,24,32,55,$ and $71$, respectively.
@@
To determine the critical temperature $T_{c}$ using GCNs, we adopt a supervised learning approach inspired by prior studies [33], but tailored to our system. Unlike traditional methods that rely on training samples labeled by a classifier with prior knowledge of $T_{c}$, we take advantage of the clear separation of phases at extreme temperatures to construct our training set. Specifically, MC spin configurations are generated at very low temperatures (e.g., $T/J=0.01$, well below the expected $T_{c}$) and very high temperatures (e.g., $T/J=8.0$, well above the expected $T_{c}$), where the system is unambiguously in the ordered and disordered phases, respectively. These configurations are labeled as “ordered” ($p_{\text{o}}=1$, $p_{\text{d}}=0$) and “disordered” ($p_{\text{o}}=0$, $p_{\text{d}}=1$) based on their temperature regimes, eliminating the need for an exact $T_{c}$ value a priori. Here $p_{\text{o}}$ and $p_{\text{d}}$ represent the classification confidences for the ordered and disordered phases, respectively. The trained GCNs then interpolates across the temperature range $[0.01,8.0]J$, predicting classification confidences $p_{\text{o}}$ and $p_{\text{d}}$ for each temperature. The intersection point where $p_{\text{o}}=p_{\text{d}}$ is taken as the estimated $T_{c}$.
@@
To verify the accuracy of the GCNs simulations, we first conducted preliminary tests on two-dimensional square lattices with periodic boundary conditions, with lattice sizes of $20\times 20$, $24\times 24$, $32\times 32$, $55\times 55$, and $71\times 71$, in order to compare with the spherical Ising model containing 500, 1000, 3000, and 5000 sites discussed later. As shown in Fig. 7, the solid lines represent the classification confidence for the ordered phase, while the dash-dot lines represent the disordered phase. As the temperature increases, the solid lines gradually decrease from 1, indicating a reduced similarity between the simulation results and the ordered phase, whereas the dash-dot lines steadily increase from 0, reflecting an increasing similarity with the disordered phase. The phase transition temperatures for different lattice sizes $L=20,24,32,55,71$ can be determined from the temperatures at which the classification confidences $p_{\text{o}}$ and $p_{\text{d}}$ intersect, yielding $T_{c}/J=2.51,2.40,2.36,2.35,2.35$, respectively. Through the simulation and analyses of different lattice scales, it is found that as the number of lattice sites increases, the phase transition temperature gradually approaches the theoretical value of $T_{c}/J=2.269$. Specifically, the values obtained are $T_{c}/J=2.51$ for the $20\times 20$ lattice and $T_{c}/J=2.35$ for the $71\times 71$ lattice. The latter demonstrates strong concordance with the predicted critical temperature. This indicates that increasing the lattice size can enhance the accuracy of phase transition temperature estimations. Moreover, the GCNs effectively captures the phase transition behavior and critical properties.
@@
After validating the accuracy of GCNs, we subsequently apply it to determine the phase transition temperature of the spherical Ising model. In numerical calculations, we select 800 temperature nodes within the range of $[0.01,8.0]J$, with an interval of $0.01J$, and perform 100 simulations for each temperature node. The results, presented in the top panel of Fig.8, indicate that the critical temperature is determined to be $T_{c}/J=2.33$, where $p_{\text{o}}=p_{\text{d}}$. This value is fairly close to those obtained from the analysis of the specific heat $C_{V}$, but it is comparatively higher than that of the square lattice, $T^{\Box}_{c}=2.269J$. Due to the inhomogeneity of the Fibonacci lattice, a complete analytical analysis is not feasible. We believe a possible reason for this is the presence of sites with 3 or 5 neighbors, which affects the spin dynamics within the system. At low temperatures, a site with a greater number of neighbors is less likely to experience a spin flip, as confirmed in Fig.4. However, as the temperature increases beyond $T/J\gtrsim 2.34$, the situation changes. Fig.4 shows that at $T/J=2.34$, the spin-flip probabilities follow the order $P_{4}>P_{5}>P_{3}$ at $T/J=2.34$, indicating that spins on sites with 3 or 5 neighbors are more resistant to flipping. This, to some extent, enhances the “stability” of the ordered phase, effectively elevating the phase transition temperature.
@@
To analyze the uncertainty of $T_{c}$, we note that it is determined by the intersection of the two curves $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$, both of which exhibit good linearity near $T_{c}$. Therefore, the uncertainty can be estimated as follows. We select an interval $T_{1},T_{2}]$ around $T_{c}$, where $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$ can be approximated as linear functions with slopes $k_{\text{o}}$ and $k_{\text{d}}$, respectively. Each point on these two curves has its own uncertainty, obtained by averaging over 100 stable spin configurations. Using error propagation, we can determine the uncertainties of $k_{\text{o}}$ and $k_{\text{d}}$, denoted as $\sigma_{\text{o}}$ and $\sigma_{\text{d}}$, respectively. Finally, the uncertainty of $T_{c}$ is given by
@@
In the bottom panel of Fig.8, we present our numerical results, yielding $T_{c}/J=2.331\pm 0.090$. Note the value $2.331J$ is obtained from the intersection of the two linear fits, which differs slightly from $T_{c}/J=2.33$, determined by the condition $p_{\text{o}}=p_{\text{d}}$.
While the GCNs machine learning method is not intended to replace traditional Monte Carlo simulations, it serves as a complementary tool to efficiently extract patterns and identify phase transitions from large configuration datasets. GCNs is particularly advantageous for automatically capturing local spin patterns on non-uniform lattices, quickly estimating critical temperatures via classification confidence and decision boundaries, and adapting to complex geometries and multi-phase systems.
III.4 Phase Transition Detection via PCA and KMeans Clustering)
Figure 9: (Left panel) The similarity of different spin configurations with respect to the two principal components. (Right panel)The cluster fraction, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature of the ferromagnetic Ising model on a spherical Fibonacci lattice for $N=1000$.
To further validate the phase transition temperatures in the spherical ferromagnetic Ising model for $N=1000$, we employed an unsupervised learning method combining Principal Component Analysis (PCA) and KMeans clustering Abdi and Williams (2010); Lancia and Scoppola (2013); Wetzel (2017b). This approach does not rely on predefined order parameters, making it well-suited for detecting statistical structure changes in complex spin systems.
We first applied PCA to the spin configurations sampled at various temperatures, each represented as a 1000-dimensional vector of ±1 values. The dimensionality was reduced to two principal components, and the resulting data were clustered using KMeans (with number of clusters set to 2). For each temperature, we calculated the fraction of samples that fall into each cluster, defined as the cluster fraction, and plotted the cluster fraction as a function of temperature.
As shown in the left panel of Fig.9, the samples are clearly grouped into two clusters, indicating that this method can effectively distinguish between the ordered and disordered phases. The right panel shows a sharp change in cluster fractions as the temperature increases, indicating a phase transition between two distinct phases. The transition point corresponds to the system’s critical temperature, estimated to be around $T_{c}/J\approx 2.32$. This result aligns with the peak of the specific heat and the critical temperature predicted by GCNs-based classification.
This unsupervised PCA plus clustering pipeline offers a powerful and intuitive visualization tool for phase transition detection, particularly in systems with curved geometry or non-uniform lattices where conventional order parameters are less accessible.
III.5 Effects of $r_{c}$ and NN
$r_{c}/R$
$N_{3}$
($w_{3}$)
$N_{4}$
($w_{4}$)
$N_{5}$
($w_{5}$)
$N_{6}$
($w_{6}$)
$N_{7}$
($w_{7}$)
$N_{8}$
($w_{8}$)
0.1298
76
(7.6%)
850
(85%)
74
(7.4%)
0.1500 -
26
(2.6%)
240
(24%)
734
(73.4%)
0.1700 - -
8
(0.8%)
722
(72.2%)
224
(22.4%)
46
(4.6%)
@@ III.5 Effects of $r_{c}$ and NN
Now, we further investigate the effects of other parameters, such as $r_{c}$ and $N$, on $T_{c}$ using Monte Carlo simulations and GCNs. First, we fix the total number of sites at $N=1000$ and vary the interaction length $r_{c}$. As $r_{c}$ increases, each spin interacts with a greater number of spins, effectively increasing the number of neighbors per site and altering the lattice structure. In this scenario, as the temperature rises, the energy cost for a spin flip becomes larger due to the higher average number of neighbors per site, leading to an effective increase in $T_{c}$. We present our numerical results in Fig.10. The top panel shows that $T_{c}/J=2.33$ for $r_{c}/R=0.1298$, $T_{c}/J=3.51$ for $r_{c}/R=0.1500$, and $T_{c}/J=3.93$ for $r_{c}/R=0.1700$, which confirms our former reasoning. Moreover, the error made by GCNs when performing phase classification can serve as an indicator for the critical behavior of a dynamical system. Consequently, the temperature at which the classification error is maximized can itself be used as a phase transition criterion Chernodub et al. (2020). To quantify this, we perform 100 independent simulations at each temperature point and compute the variance of the classification confidence. As shown in the middle panel, the maximum classification errors occur at $T_{c}/J=2.34$ for $r_{c}/R=0.1298$, $T_{c}/J=3.53$ for $r_{c}/R=0.1500$, and $T_{c}/J=3.95$ for $r_{c}/R=0.1700$, which agree well with the phase transition temperature obtained from the intersection of the two phases ($p_{\text{o}}$, $p_{\text{d}}$) in classification. This further supports the reliability of the maximum confusion criterion. The bottom panel lists the fraction of sites with different numbers of neighbors for various $r_{c}$ values. For example, when $r_{c}=0.1500R$, $73.4\%$ of the sites have six neighbours, indicating that most of the lattice resembles a triangular structure. Interestingly, the estimated $T_{c}/J=3.51$ is relatively close to the exact value $T_{c}/J=4/\text{ln}3\approx 3.64$ of the planar triangular lattice as reported in Ref.Baxter (2016); Wannier (1950, 1973); Houtappel (1950a, b). Since this model also includes some four-neighbor and five-neighbor sites, its spin flip energy is lower than that of the six-neighbor sites, which results in a relatively lower $T_{c}$. For $r_{c}/R=0.1700$, although a similar fraction of sites have six neighbors ($w_{6}=72.2\%$), the number of sites with more neighbors increases significantly ($w_{7}=22.4\%$ and $w_{8}=4.6\%$). This enhanced “neighbor interaction” effectively strengthens spin correlations, leading to a higher phase transition temperature of $T_{c}/J=3.93$.
NN $r_{c}/R$ $N_{3}$ ($w_{3}$) $N_{4}$ ($w_{4}$) $N_{5}$ ($w_{5}$)
500 0.1813 68 (11.6%) 418 (83.6%) 14(2.8%)
1000 0.1298 76 (7.6%) 850 (85%) 74 (7.4%)
3000 0.0747 186 (6.2%) 2740 (91.3%) 74 (2.5%)
5000 0.0582 112 (2.2%) 4740 (94.8%) 148 (3%)
@@
Figure 11: (Top panel) Classification confidences, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature for $(N,r_{c}/R)=(1000,0.1298)$, (3000,0.0747) and (5000,0.0582). (Bottom panel) Corresponding fractions of sites with different numbers of neighbors.
@@
Next, we vary the number of lattice sites and adjust $r_{c}$ accordingly to ensure the majority of sites have four neighbors. However, due to the absence of translational and rotational symmetry in the spherical Fibonacci lattice, this adjustment does not necessarily guarantee that lattices with different site numbers share similar geometric structures. This contrasts with the 2D square lattice, where structural consistency is maintained. In addition to the previously discussed case of $N=1000$, $r_{c}/R=0.1298$, we also select $N=500$, $r_{c}/R=0.1813$; $N=3000$, $r_{c}/R=0.0747$ and $N=5000$, $r_{c}/R=0.0582$, ensuring that the fractions of sites with four neighbors in the latter three cases are 83.6%, 91.3% and 94.8%, respectively. Interestingly, this suggests that the “irregular area” gradually diminishes as the effective curvature of the spherical lattice decreases. The numerical results are visualized in the top panel of Fig.11. Similarly, we determine the phase transition temperatures using two different methods mentioned before. First, by identifying the intersection of the classification confidences, we obtain $T_{c}/J=2.33$, $2.33$, $2.34$ and $2.35$ for $N=500$, $1000$, $3000$ and $5000$, respectively. Second, using linear fitting near $T_{c}$ (see the bottom panel of Fig.8), the corresponding phase transition temperatures are giving $T_{c}/J=2.331\pm 0.0742$, $2.331\pm 0.0900$, $2.348\pm 0.0615$ and $2.358\pm 0.0496$ for the corresponding system sizes. Compared with the results for planar square lattice with a similar number of sites and periodic boundary conditions, as the fraction of sites with four neighbors increases, the $T_{c}$ of the spherical Ising model gradually approaches that of the planar square lattice. All these findings suggest that even for a simple system like the Ising model, the spherical Fibonacci lattice exhibits properties distinct from those of planar lattices, similar to the observations in Ref.Pochinok et al. (2023).